Find the slope, if it exists, of the line containing the pair of points.
(1,10) and (8,-9)
the slope is:

Answer by fcabanski(1391)

Slope is (y2-y1)/(x2-x1).


A point is (x,y) and it doesn't matter which point you call 1 and which you call 2.


Call (1,10) point 1 and (8,-9) point 2.
(-9-10)/(8-1) = -19/7 is the slope.


Call (8,-9) point 1 and (1,10) point 2.
(10-(-9))/(1-8) = 19/-7 = -19/7 is the slope. It's the same.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com